OpenStudy (anonymous):

What is the slope of the line that contains the points (-1,-1) and (3,15)

OpenStudy (aroub):

15+1/3+1=16/4=4 the slope is 4

OpenStudy (anonymous):

use the formula

OpenStudy (anonymous):

\[change \in y\]---------\[change \in x\]

OpenStudy (anonymous):

-1-15 / -1-3 -16/-4 gradient=4

OpenStudy (anonymous):

nice